solve system using addition method
-6x+y=6
-24x-4y=24
solution
(?,?)

Answer by checkley77(12844)

-6x+y=6 multiply by 4 & add
-24x-4y=24
-24x+4y=24
--------------------
-48x+0y=48
x=48/-48
x=-1 ans.
-6*-1+y=6
6+y=6
y=6-6
y=0 ans.